Alex buys 36 cans of vegetables. He pays 80 cents for tomatoes and 40 cents for a can of corn, for a total cost of $20.80. How many cans of tomatoes did he buy?
Answer by scott8148(6628)

"36 cans of vegetables" ___ c + t = 36 ___ c = 36 - t

"80 cents for tomatoes and 40 cents for a can of corn, for a total cost of $20.80" ___ 80t + 40c = 2080

substituting ___ 80t + 40(36 - t) = 2080

80t + 1440 - 40t = 2080

40t = 640 ___ t = 16